[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Jun 27, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Jun 27 15:15:19 UTC 2017


Here is today's summary of NHC computer operations:
--- Matt Sardi

----------

TAFB:
1. NCO Windows Sysadmin Marc Laurent determined that the noise from the  
Situational Awareness PC was caused by a failing graphics card. A new card  
has been shipped overnight and will be installed soon.
HSU:
1. Interactive cone graphic did not update properly for Dora advisory 8  
(21z). Dave Z. was contacted and fixed the issue. Should work for  
subsequent advisories.

2. 18Z ECMWF SHIPS intensity guidance did not come in despite having been  
pulled several times between 1505Z-1615Z.

Reminder, getting ECMWF SHIPS to run requires the forecaster to resubmit  
guidance after the 06/18Z cycle ECMWF data become available.
Other items:
1. NCO will be performing a bandwidth upgrade on Wednesday, 6/28; network  
service degradation/interruptions are possible during two 20 minute windows  
(1530-1550Z & 1620-1640Z). TAFB/HSU service backup may be required if  
outages last longer than expected. Work was originally to be started on  
Tuesday, June 20 at 1530Z; postponed to Wednesday, June 28 at 1500Z (11 am  
EDT).

2. The GOES-16 data outage in N-AWIPS/AWIPS has been resolved;
There are number of new derived channels being ingested into AWIPS2 D2D  
after we reconfigured the data ingest from the GRB servers. More info soon.

3. Doug Gaer (NCO) and Chris Mello (TSB) are installing AWIPS Build 17.1.1.  
on the NHCR testbed system today. This will not affect operations.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170627/9b70286d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list